OpenStudy (anonymous):

Translate this word problem as a system of equations and then solve using substitution. The sum of two numbers is 84. One number is three times the other. Find the numbers. Put your answer in this form: The numbers are 3 and 6.

OpenStudy (anonymous):

need helpfast

OpenStudy (anonymous):

need the answer fast

OpenStudy (anonymous):

X+y=84 y=3x 4x=84 X=21 Y=84-21=63